### v3

@@ D.1 The Case of $\mu_{f}>0$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{1}{T+1+2\kappa_{f}}\left(\frac{1}{T-t+1}+\frac{1}{t+2\kappa_{f}}\right)$
$\leq$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{1+\log T+\frac{1}{1+2\kappa_{f}}+\log\frac{T+2\kappa_{f}}{1+2\kappa_{f}}}{T+1+2\kappa_{f}}$
$\leq$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{4(M^{2}+\sigma^{2})(1+\log T)}{\mu_{f}(T+2\kappa_{f})}=O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (38)
@@
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+1+4\kappa_{f}}{s-1+4\kappa_{f}}=\frac{(t+4\kappa_{f})(t+1+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}.$
Thus, we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{(1-\frac{1}{1+2\kappa_{f}})D_{\psi}(x,x^{1})}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\frac{1}{\mu_{f}^{2}(1+2\kappa_{f})^{2}}}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\cdot\frac{4}{\mu_{f}^{2}(t+1+4\kappa_{f})^{2}}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\cdot\frac{2}{\mu_{f}(t+1+4\kappa_{f})})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}\right$.
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+\sum_{t=2}^{T}\frac{4}{(\Gamma_{T}/\Gamma_{t-1}-1)(t-1+4\kappa_{f})(t+1+4\kappa_{f})}\right]$
$=$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}$
$+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{2(1+4\kappa_{f})}{(1+2\kappa_{f})T(T+1+8\kappa_{f})}+\sum_{t=2}^{T}\frac{4(t+4\kappa_{f})}{(T+1-t)(T+t+8\kappa_{f})(t+1+4\kappa_{f})}\right]$
$=$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{4(t+4\kappa_{f})}{(T+1-t)(T+t+8\kappa_{f})(t+1+4\kappa_{f})}$
$\leq$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{4}{2T+1+8\kappa_{f}}\left(\frac{1}{T+1-t}+\frac{1}{T+t+8\kappa_{f}}\right)$
$\leq$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{4(1+\log T+\log\frac{2T+8\kappa_{f}}{T+8\kappa_{f}})}{2T+1+8\kappa_{f}}$
$\leq$ $\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{8(M^{2}+\sigma^{2})(1+\log 2T)}{\mu_{f}(2T+1+8\kappa_{f})}=O\left(\frac{L(1+\kappa_{f})D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (39)
@@
$\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}.$
@@
Note that we have $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
@@
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\begin{cases}\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{t-1}&t\leq\tau\\ \left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(t-\tau+1+4\kappa_{f})(t-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}&t\geq\tau+1\end{cases}.$
So we know
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\eta_{1}^{2}}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}$
$+2\mu_{f}(M^{2}+\sigma^{2})\left[\underbrace{\sum_{t=2}^{\tau}\frac{\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}}_{\text{I}}+\underbrace{\sum_{t=\tau+1}^{T}\frac{\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}}_{\text{II}}\right].$ (40)
Note that we can bound
$\frac{1}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}=$ $\frac{1}{\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{2(1+4\kappa_{f})}-2\kappa_{f}}$
$\overset{(a)}{\leq}$ $\frac{1}{\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}(1+2\kappa_{f})-2\kappa_{f}}=\frac{1}{2\kappa_{f}\left[\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}-1\right]+\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}}$
$\overset{(b)}{\leq}$ $\left(1-\frac{1}{\eta+2\kappa_{f}}\right)^{\tau-1}\leq\exp\left(-\frac{\tau-1}{\eta+2\kappa_{f}}\right)=\exp\left(\frac{-\tau}{\eta+2\kappa_{f}}+\frac{1}{\eta+2\kappa_{f}}\right)$
$\overset{(c)}{\leq}$ $\exp\left(-\frac{T}{2(\eta+2\kappa_{f})}+1\right),$ (41)
where $(a)$ holds due to $T-\tau\geq 0$, $(b)$ is by $\kappa_{f}\geq 0$ and $\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\geq 1$, $(c)$ is from $\tau\geq\frac{T}{2}$, $\eta+\kappa_{f}>1$ and $\kappa_{f}\geq 0$. We can also bound
$\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}$
$=$ $\frac{1}{(1+2\kappa_{f})\left[\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-1}\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{2(1+4\kappa_{f})}-2\kappa_{f}\right]}$
$\leq$ $\frac{1}{(1+2\kappa_{f})\left[\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{2(1+4\kappa_{f})}-2\kappa_{f}\right]}=\frac{2(1+4\kappa_{f})}{(1+2\kappa_{f})(T-\tau+1)(T-\tau+2+8\kappa_{f})}$
$\overset{(d)}{\leq}$ $\frac{2(1+4\kappa_{f})}{(1+2\kappa_{f})(T-\frac{T+1}{2}+1)(T-\frac{T+1}{2}+2+8\kappa_{f})}=\frac{8(1+4\kappa_{f})}{(1+2\kappa_{f})(T+1)(T+3+16\kappa_{f})}$
$\overset{(e)}{\leq}$ $\frac{8}{T+3+16\kappa_{f}},$ (42)
where $(d)$ is by $\tau\leq\frac{T+1}{2}$ and $(e)$ is by $T\geq 1$. Besides, there is
$\text{I}=$ $\frac{1}{\mu_{f}^{2}(\eta+2\kappa_{f})(\eta+2\kappa_{f}-1)}\sum_{t=2}^{\tau}\frac{1}{\left(\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\right)^{\tau-t+1}\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}-1}$
$\overset{(f)}{\leq}$ $\frac{1}{\mu_{f}^{2}(\eta+2\kappa_{f})(\eta+2\kappa_{f}-1)\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}}\sum_{t=2}^{\tau}\frac{1}{\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(1+4\kappa_{f})(2+4\kappa_{f})}-1}$
$=$ $\frac{1}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\sum_{t=2}^{\tau}\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{(T-\tau)(T-\tau+3+8\kappa_{f})}=\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\cdot\frac{\tau-1}{(T-\tau)(T-\tau+3+8\kappa_{f})}$
$\overset{(g)}{\leq}$ $\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\cdot\frac{\frac{T+1}{2}-1}{(T-\frac{T+1}{2})(T-\frac{T+1}{2}+3+8\kappa_{f})}$
$=$ $\frac{(1+4\kappa_{f})(2+4\kappa_{f})}{\mu_{f}^{2}(\eta+2\kappa_{f})^{2}}\cdot\frac{2}{T+5+16\kappa_{f}}\overset{(h)}{\leq}\frac{32}{\mu_{f}^{2}(T+5+16\kappa_{f})},$ (43)
where $(f)$ is by $\tau-t\geq 0$ (w.l.o.g., we can assume $\tau\geq 2$, otherwise, there is $\text{I}=0\leq\frac{8}{\mu_{f}^{2}(T+5+16\kappa_{f})}$) and $\frac{\eta+2\kappa_{f}}{\eta+2\kappa_{f}-1}\geq 1$, $(g)$ is due to $\tau\leq\frac{T+1}{2}$ and $(h)$ is by $\eta+\kappa_{f}>1$. We also have
$\text{II}=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=\tau+1}^{T}\frac{1}{\left[\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(t-\tau+4\kappa_{f})(t-\tau+1+4\kappa_{f})}-1\right](t-\tau+2+4\kappa_{f})(t-\tau+4\kappa_{f})}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{\left[\frac{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})}{(t+4\kappa_{f})(t+1+4\kappa_{f})}-1\right](t+2+4\kappa_{f})(t+4\kappa_{f})}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{t+1+4\kappa_{f}}{\left[(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})-(t+4\kappa_{f})(t+1+4\kappa_{f})\right](t+2+4\kappa_{f})}$
$\leq$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1+4\kappa_{f})(T-\tau+2+4\kappa_{f})-(t+4\kappa_{f})(t+1+4\kappa_{f})}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1-t)(T-\tau+2+8\kappa_{f}+t)}$
$=$ $\frac{4}{\mu_{f}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{2T-2\tau+3+8\kappa_{f}}\left(\frac{1}{T-\tau+1-t}+\frac{1}{T-\tau+2+8\kappa_{f}+t}\right)$
$\leq$ $\frac{4(1+\log(T-\tau)+\log\frac{2T-2\tau+2+8\kappa_{f}}{T-\tau+2+8\kappa_{f}})}{\mu_{f}^{2}(2T-2\tau+3+8\kappa_{f})}\leq\frac{4(1+\log T)}{\mu_{f}^{2}(T+2+8\kappa_{f})},$ (44)
where we use $\frac{T}{2}\leq\tau\leq\frac{T+1}{2}$ in the last inequality.
Plugging (41), (42), (43) and (44) into (40), we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $2LD_{\psi}(x,x^{1})\exp\left(-\frac{T}{2(\eta+2\kappa_{f})}+1\right)+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\cdot\frac{8}{T+3+16\kappa_{f}}$
$+2\mu_{f}(M^{2}+\sigma^{2})\left[\frac{32}{\mu_{f}^{2}(T+5+16\kappa_{f})}+\frac{4(1+\log T)}{\mu_{f}^{2}(T+2+8\kappa_{f})}\right]$
$=$ $O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2\eta+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (45)
Theorem D.2.
@@
When $\mu_{f}>0$ and $\mu_{h}=0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ in Lemma 4.3 holds, we will have with probability at least $1-\delta$
$F(x^{T+1})-F(x)\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left[\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right].$ (46)
Now observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\eta_{t}\Gamma_{t}=\begin{cases}\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{f}}&t\geq 2\\ \eta_{1}&t=1\end{cases},$
where $\Gamma_{t\in\left[T\right]}\coloneqq\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}$. Hence, (46) can be rewritten as
$F(x^{T+1})-F(x)\leq 2\left(1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}\right)\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$ (47)
Now let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ for our three choices respectively:
@@
First, we consider $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$. We can find $1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}=1+\frac{1}{1-\mu_{f}\eta_{2}}\leq 3$, $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+2\kappa_{f}}{s-1+2\kappa_{f}}=\frac{t+2\kappa_{f}}{1+2\kappa_{f}}.$
Hence, using (47), we have
$F(x^{T+1})-F(x)\leq 6\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
@@
Next, for the case of $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, there are $1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}=1+\frac{1}{1-\mu_{f}\eta_{2}}\leq 4$, $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+1+4\kappa_{f}}{s-1+4\kappa_{f}}=\frac{(t+4\kappa_{f})(t+1+4\kappa_{f})}{2(1+4\kappa_{f})(1+2\kappa_{f})}.$
Thus, we have
$F(x^{T+1})-F(x)\leq 8\left(\frac{D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$

### v4

@@ D.1 The Case of $\mu_{f}>0$
$=\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}(T+1+2\kappa_{f})}\sum_{t=1}^{T}\frac{1}{T-t+1}+\frac{1}{t+2\kappa_{f}}$
$\leq\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}(T+1+2\kappa_{f})}\left(1+\log T+\frac{1}{1+2\kappa_{f}}+\log\frac{T+2\kappa_{f}}{1+2\kappa_{f}}\right)$
$\leq\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{4(M^{2}+\sigma^{2})(1+\log T)}{\mu_{f}(T+1+2\kappa_{f})}$
$=O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (37)
@@
$\gamma_{t\in\left[T\right]}$ $=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})\prod_{s=2}^{t}(1-\frac{2}{s+1+4\kappa_{f}})}$
$=\frac{t+4\kappa_{f}}{\mu_{f}(1+2\kappa_{f})(1+4\kappa_{f})}=\frac{t+4\kappa_{f}}{1+4\kappa_{f}}\eta_{1}.$ (38)
Hence, using (35) and (38), we have
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{(1+4\kappa_{f})(\eta_{1}^{-1}-\mu_{f})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}t+4\kappa_{f}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{(t+4\kappa_{f})\eta_{t}}{\sum_{s=t}^{T}s+4\kappa_{f}}$
$=\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{8(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{t+4\kappa_{f}}{(t+1+4\kappa_{f})(T-t+1)(T+t+8\kappa_{f})}$
$\leq\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{8(M^{2}+\sigma^{2})}{\mu_{f}}\sum_{t=1}^{T}\frac{1}{(T-t+1)(T+t+8\kappa_{f})}$
$=\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{8(M^{2}+\sigma^{2})}{\mu_{f}(2T+1+8\kappa_{f})}\sum_{t=1}^{T}\frac{1}{T-t+1}+\frac{1}{T+t+8\kappa_{f}}$
$\leq\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{8(M^{2}+\sigma^{2})}{\mu_{f}(2T+1+8\kappa_{f})}\left(1+\log T+\log\frac{2T+8\kappa_{f}}{T+8\kappa_{f}}\right)$
$\leq\frac{4(1+4\kappa_{f})LD_{\psi}(x,x^{1})}{T(T+1+8\kappa_{f})}+\frac{8(M^{2}+\sigma^{2})(1+\log 2T)}{\mu_{f}(2T+1+8\kappa_{f})}$
$=O\left(\frac{L(1+\kappa_{f})D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (39)
@@
$\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases},$
@@
where $\tau=\left\lceil\frac{T}{2}\right\rceil$. Note that we have $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
@@
$\gamma_{t\in\left[T\right]}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}=\begin{cases}\frac{\eta_{1}}{(1-\mu_{f}\eta_{1})^{t-1}}&t\leq\tau\\ \frac{\eta_{1}(t-\tau+1+4\kappa_{f})}{(1-\mu_{f}\eta_{1})^{\tau-1}(1+4\kappa_{f})}&t\geq\tau+1\end{cases}.$ (40)
So, there is
[…]
$\leq\frac{(1-\mu_{f}\eta_{1})^{\tau}}{\eta_{1}}\leq(\eta_{1}^{-1}-\mu_{f})\exp\left(-\mu_{f}\eta_{1}(\tau-1)\right)$
$=2L\exp\left(-\frac{\tau-1}{1+2\kappa_{f}}\right)\leq 2eL\exp\left(-\frac{T}{2+4\kappa_{f}}\right).$ (41)
Now, we observe that
$\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}=\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}+\sum_{t=\tau+1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}=\underbrace{\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta_{1}}{\sum_{s=t}^{T}\gamma_{s}}}_{\mathrm{I}}+\underbrace{\sum_{t=\tau+1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}}_{\mathrm{II}}.$
We bound terms $\mathrm{I}$ and $\mathrm{II}$ separately in the following.
  • Term $\mathrm{I}$: When $t\leq\tau$, we have
[…]
    implying that
    $\mathrm{I}$ $\leq\frac{2\tau\eta_{1}(1+4\kappa_{f})}{(T-\tau+1)(T-\tau+2+8\kappa_{f})}\overset{\tau\leq\frac{T+1}{2}}{\leq}\frac{4\eta_{1}(1+4\kappa_{f})}{T+3+16\kappa_{f}}$
    $=\frac{4(1+4\kappa_{f})}{\mu_{f}(1+2\kappa_{f})(T+3+16\kappa_{f})}\leq\frac{8}{\mu_{f}(T+3+16\kappa_{f})}.$
  • Term $\mathrm{II}$: When $t\geq\tau+1$, we have
[…]
    implying that
    $\mathrm{II}$ $\leq\sum_{t=\tau+1}^{T}\frac{4}{\mu_{f}(T-t+1)(T+t-2\tau+2+8\kappa_{f})}$
    $=\frac{4}{\mu_{f}(2T-2\tau+3+8\kappa_{f})}\sum_{t=\tau+1}^{T}\frac{1}{T-t+1}+\frac{1}{T+t-2\tau+2+8\kappa_{f}}$
    $\leq\frac{4}{\mu_{f}(2T-2\tau+3+8\kappa_{f})}\left(1+\log(T-\tau)+\log\frac{2T-2\tau+2+8\kappa_{f}}{T-\tau+2+8\kappa_{f}}\right)$
    $\overset{\frac{T}{2}\leq\tau\leq\frac{T+1}{2}}{\leq}\frac{4(1+\log T)}{\mu_{f}(T+2+8\kappa_{f})}.$
Therefore, we obtain
$\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\leq\frac{8}{\mu_{f}(T+3+16\kappa_{f})}+\frac{4(1+\log T)}{\mu_{f}(T+2+8\kappa_{f})}\leq\frac{12(1+\log T)}{\mu_{f}(T+2+8\kappa_{f})}.$ (42)
We combine (35), (41), and (42) to conclude
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{2eLD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2+4\kappa_{f}}\right)}+\frac{24(M^{2}+\sigma^{2})(1+\log T)}{\mu_{f}(T+2+8\kappa_{f})}=O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$ (43)
Theorem D.2 (Full version of Theorem 3.7).
@@
When $\mu_{f}>0$ and $\mu_{h}=0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ in Lemma 4.3 holds, we have, with probability at least $1-\delta$,
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}$. Now, let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ for the three choices respectively:
@@
First, we consider $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$. We can find $1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}=1+\frac{1}{1-\mu_{f}\eta_{2}}\leq 3$. Hence, using (44), we have
$F(x^{T+1})-F(x)\leq 6\left(\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$
@@
Next, for the case of $\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}$, there is $1+\max_{2\leq t\leq T}\frac{1}{1-\mu_{f}\eta_{t}}=1+\frac{1}{1-\mu_{f}\eta_{2}}\leq 4$. Thus, using (44), we have
$F(x^{T+1})-F(x)\leq 8\left(\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\right).$